A cylindrical metal has a height of 0.27 m and a radius of 0.11 m. The electric field is directed outward along the entire surface of the can (including the top and bottom) with uniform and magnitude of 4.0 x 10 5N/C. How much charge does the can contain?


1
Expert's answer
2021-03-31T16:14:39-0400
"A=2\pi r^2+2\pi rh=2\pi r(r+h)\\EA=\frac{q}{\epsilon_0}\\\\2\pi r(r+h)E=\frac{q}{\epsilon_0}
\\2\pi (0.11)(0.11+0.27)(4\cdot10^5)=\frac{q}{8.85\cdot10^{-12}}\\q=9.3\cdot10^{-7}\ C"